Tia is 41 years old and her daughter is 5 years old. After x years, her daughter age will be half her age. What is the value of x
Answer by ikleyn(52873)

.
Tia is 41 years old and her daughter is 5 years old.
After x years, her daughter age will be half her age. What is the value of x
~~~~~~~~~~~~~~~

Write equation as you red the problem

    41 + x = 2*(5+x).


Now simplify and find x

    41 + x = 10 + 2x

    41 - 10 = 2x - x

       31   =    x


ANSWER.  It will happen in 31 years.

Solved and explained.

--------------

Notice that the daughter was born when the mother was 41-5 = 36 years old.

So it is clear without calculations that the next event will happen when they will be 36 and 72 years old,
respectively, which will happen 36 - 5 = 31 years from now.